Please answer all parts of the following question (you must show all work):

1. Lumyn Electronics is a family owned business. The company sold 128 TV remote controls in July and incurred a total cost of $2,800. In November they sold 248 TV remote controls and incurred a total cost of $4,600. Based on this information, find the following:

a) The marginal cost and monthly fixed cost for this company.

b) A linear function that expresses the company's cost.

c) The revenue and profit (or loss) for this company for the months of July and November, if the average price for each TV remote control sold was $20.
Answer #1

Hi! Welcome to Chegg!

1.

a.

Let the cost function be:

C = A + BQ

where,

A = Monthly fixed cost

B = Marginal cost

Q = Sales

So, we have:

July data: C = 2800 = A + B(128)

November data: C = 4600 = A + B(248)

Solving for A and B, we get:

2800 - 128B = 4600 - 248B

1800 = 120B

Marginal cost = B = 15

Fixed cost = A = 2800 - 128(15) = 880.

b.

Linear cost function:

C = 880 + 15Q.

c.

Price = $20

Profit in July = PQ - C = 20(128) - 2800 = - 240 (Loss)

Profit in November = PQ - C = 20(248) - 4600 = 360.

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!